# Higher-dimensional Teter rings via the canonical trace

### v1

@@ 2.4. Numerical semigroup rings
$\operatorname{PF}(H):=\{h\in H\,|\,\alpha+h\in H\,\;{\it for\;all}\;\,h\in H\},$
$\operatorname{Ap}(H,a_{1})=\{h\in H\mid h-a_{1}\in H\}.$
@@
and we have $\operatorname{r}(R_{H})=\#(\operatorname{PF}(H))=:\operatorname{type}(H)$. Next, we describe several classes of numerical semigroups that can be characterized in terms of $\operatorname{PF}(H)$.
Definition 2.16.
Let $H$ be a numerical semigroup.
@@
    Suppose that $\operatorname{F}(H)$ is even. We say that $H$ is pseudo-symmetric if $\operatorname{PF}(H)=\{\frac{\operatorname{F}(H)}{2},\operatorname{F}(H)\}$.
@@
$\Omega_{H}^{+}=\{m\in\mathbb{N}\mid t^{m}\in\omega_{R_{H}}\},$
@@
$\Omega_{H}^{-}=\{m\in\mathbb{N}\mid m-\alpha_{i}\in H{\it\;\;for\;all\;\;}i\in[n]\}.$
According to [herzog2019trace, Lemma 1.1], the canonical trace ideal $\operatorname{tr}_{R_{H}}(\omega_{R_{H}})$ of $R_{H}$ is obtained by
$\operatorname{tr}_{R_{H}}(\omega_{R_{H}})=\omega_{R_{H}}\omega_{R_{H}}^{-1}.$
Thus, the set $\operatorname{tr}_{R_{H}}(H)$ corresponding to the generators of $\operatorname{tr}_{R_{H}}(\omega_{R_{H}})$ is given by
$\operatorname{tr}(H)=\Omega_{H}^{+}+\Omega_{H}^{-}.$
@@ 3. Graded Teter rings and canonical traces
    Assume that $\dim R=0$ and $R$ is not Gorenstein, then the following are equivalent:
@@
$\varphi(\omega_{R})=\mathfrak{m}\quad\text{or}\quad\varphi\text{ is injective and }\operatorname{embdim}\!\bigl(R/\varphi(\omega_{R})\bigr)\leq\dim R.$
@@
(1) $\Rightarrow$ (4): Choose $\varphi:\omega_{R}\to R$ as the composite of an isomorphism $\omega_{R}\cong I$ with the inclusion $I\hookrightarrow R$. Then $\operatorname{embdim}\!\bigl(R/\varphi(\omega_{R})\bigr)=\operatorname{embdim}(R/I)\leq\dim R$, and hence $R$ is Teter.
(4) $\Rightarrow$ (1): By (a) and definition, there exists a graded $R$-homomorphism $\varphi:\omega_{R}\to R$ such that $\varphi$ is injective and $\operatorname{embdim}\bigl(R/\varphi(\omega_{R})\bigr)\leq\dim R$. Set $I=\varphi(\omega_{R})$. As $R$ is not Gorenstein, we have $I\neq R$; consequently $\omega_{R}\cong I$ and $\operatorname{embdim}(R/I)\leq\dim R$, establishing (1).
@@
When $\dim R=0$, it follows from Section  3 (a) that $R$ is Teter in the sense of [Teter1974SocleFactor]. In this case, [elias2017teter, Theorem 3.4] yields $r(R)=\operatorname{embdim}(R)=\text{codim}(R)$. Hence it suffices to treat the case $\dim R>0$. By Section  3 (b), the localization $R_{{\mathfrak{m}}}$ is a local Teter ring. Therefore, by [puthenpurakal2025higher, Proposition 5.15] we have $r(R_{{\mathfrak{m}}})=\operatorname{embdim}(R_{{\mathfrak{m}}})-\dim(R_{{\mathfrak{m}}})$. Since $r(R_{{\mathfrak{m}}})=r(R)$, $\operatorname{embdim}(R_{{\mathfrak{m}}})=\operatorname{embdim}(R)$, $\dim(R_{{\mathfrak{m}}})=\dim(R)$ (see [bruns1998cohen, Proposition 1.5.15]), we conclude that $r(R)=\operatorname{embdim}(R)-\dim(R)=\text{codim}(R)$. ∎
Recall that $r_{0}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$.
@@
Assume that $\omega_{R}]_{-a_{R}}$ contains an $R$-torsion free element of $\omega_{R}$. If $\operatorname{tr}(\omega_{R})]_{\text{indeg}_{R}({\mathfrak{m}})}$ contains a nonzero-divisor, then there is a graded ideal $I$ of $R$ such that $\omega_{R}\cong I$ as a graded $R$-module and
$r_{0}(R)\leq\operatorname{embdim}(R)-\operatorname{embdim}(R/I).$
@@
By § 2.3, there exists a graded ideal $I\subsetneq R$ such that $\omega_{R}\cong I$. Then, $\operatorname{embdim}(R/I)\leq\operatorname{embdim}(R)-r_{0}(R)$. This yields the desired inequality. ∎
@@
Assume that $r_{0}(R)\geq\text{codim}(R)$, $\omega_{R}]_{-a_{R}}$ contains an $R$-torsion free element of $\omega_{R}$, and that $\operatorname{tr}(\omega_{R})]_{\text{indeg}_{R}({\mathfrak{m}})}$ contains a nonzero-divisor, then $R$ is Teter.
@@
$\operatorname{embdim}(R)-\dim(R)=\text{codim}(R)\leq r_{0}(R)\leq\operatorname{embdim}(R)-\operatorname{embdim}(R/I).$
@@
Since $R$ is Teter, by Section  3 there exists a graded ideal $I\subset R$ such that $\omega_{R}\cong I$ and $\dim_{R_{0}}(I_{1})\geq\text{codim}(R)$. In this case $I_{1}\neq 0$, and since $r_{0}(R)=\dim_{R_{0}}(I_{1})\geq\text{codim}(R)$, we have $r_{0}(R)\geq\text{codim}(R)$. Therefore, by Section  3 we obtain $\text{codim}(R)\leq r_{0}(R)\leq r(R)=\text{codim}(R)$. Hence $r(R)=r_{0}(R)$, so $R$ is level. Because $R$ is Teter with $\dim(R)>0$, it is in particular generically Gorenstein by Section  3 (c). Since $R$ has positive Krull dimension, is level, and is generically Gorenstein, the same argument as in the proof of [bruns1998cohen, Theorem 4.4.9] shows that $I_{1}\ (\cong[\omega_{R}]_{-a_{R}})$ contains a non-zerodivisor of $R$. Consequently $\operatorname{tr}_{R}(\omega_{R})]_{1}=[\operatorname{tr}_{R}(I)]_{1}$ contains $I_{1}$ and, in particular, contains a non-zerodivisor of $R$. ∎
@@
Let $R$ be a Cohen–Macaulay standard graded Teter ring. By Section  3, if $\dim(R)>0$ then $R$ is level, so it suffices to consider the case $\dim(R)=0$. Although the fact that an Artinian graded Teter ring is level is probably well known to experts, we include a proof for completeness. In this case there exists a standard graded Artin Gorenstein ring $G$ such that $R\cong G/\operatorname{Soc}(G)$ and $\operatorname{embdim}(R)=\operatorname{embdim}(G)$, where $\operatorname{Soc}(G):=\{a\in G:a{\mathfrak{m}}_{G}=0\}$. Let $s$ be the socle degree of $G$ and put $n=\operatorname{embdim}(R)$. Then the $h$-vectors of $G$ and $R$ can be written as $h(G)=(h_{0},\dots,h_{s(G)})$ and $h(R)=(h_{0},\dots,h_{s(G)-1})$. Since $G$ is Gorenstein, its $h$-vector is symmetric (see [stanley1978hilbert, Theorem 4.1]), hence $h_{s(G)-1}=h_{1}=\dim_{R_{0}}R_{1}=n$. Therefore $R$ is level by [elias2017teter, Theorem 3.12]. ∎
@@
    Section  3 fails in general in the semi-standard graded setting, and there exist examples in which the difference between the Cohen–Macaulay types $r(R)$ and $r_{0}(R)$ can take any prescribed value (see Section  4).
@@
    In Section  3 (c), (4) $\Rightarrow$ (3) does not hold in general. Let $\mathbb{k}$ be a field and let $A=\mathbb{k}[x,y]/(x^{3},y^{4},xy)$. Note that $B=A_{{\mathfrak{m}}}$ is an Artinian Teter local ring in the sense of [Teter1974SocleFactor] (see [gasanova2022rings, Example 4.3 (2)]). Set $R=A[z]$ and regard $R$ as a standard graded ring by declaring $\operatorname{deg}(x)=\operatorname{deg}(y)=\operatorname{deg}(z)=1$. By [puthenpurakal2025higher, Example 5.15], we have $R_{{\mathfrak{m}}_{R}}\cong B[z]_{{\mathfrak{m}}_{B}B[z]+zB[z]}$ is a Teter local ring in the sense of [puthenpurakal2025higher]. However, since $R$ is not level, $R$ is not Teter by Section  3.
The following statement is proved (for standard graded rings) in the course of the proof of [bruns1998cohen, Theorem 4.4.9], but the same argument goes through unchanged without the standard graded hypothesis. For completeness, we include a proof.
Remark 3.12 (cf. [bruns1998cohen, Theorem 4.4.9]).
@@
We may assume that $R_{0}$ is an infinite field. Put b=a(R)b=-a(R) and $V=[\omega_{R}]_{b}$. Choose a graded surjection $G\twoheadrightarrow\omega_{R}^{\vee}$ with $G$ a finite free graded module, and dualize:
$\omega_{R}\xrightarrow{\ \varphi\ }\omega_{R}^{\vee\vee}\hookrightarrow G^{\vee}=:F.$
Since $R$ is generically Gorenstein, $\varphi$ is injective; hence we view $\omega_{R}$ as a graded submodule of the free module $F$. For each $\mathfrak{p}\in\operatorname{Ass}R$ the localization $(\omega_{R})_{\mathfrak{p}}\cong R_{\mathfrak{p}}$ is free of rank $1$. If $V\subset(\mathfrak{p}F)_{b}$ held for some $\mathfrak{p}$, the $R$-submodule generated by $V$ would give $\omega_{R}\subset\mathfrak{p}F$, and localizing at $\mathfrak{p}$ would force $(\omega_{R})_{\mathfrak{p}}\subset\mathfrak{p}R_{\mathfrak{p}}F_{\mathfrak{p}}$, which is impossible because the left-hand side is free of rank $1$. Hence $(\mathfrak{p}F)_{b}$ is a proper $R_{0}$-subspace of $V$ for every $\mathfrak{p}\in\operatorname{Ass}R$. As $R_{0}$ is infinite, a finite union of proper subspaces cannot cover $V$. We may therefore choose $x\in V\setminus\bigcup_{\mathfrak{p}\in\operatorname{Ass}R}(\mathfrak{p}F)_{b}$. Consider the graded homomorphism $\mu:R\to\omega_{R}(b)$, $1\mapsto x$. For each $\mathfrak{p}\in\operatorname{Ass}R$ the localization $\mu_{\mathfrak{p}}$ is nonzero, because $x\notin\mathfrak{p}F$.
If $\operatorname{ker}(\mu)\neq 0$, let $\mathfrak{q}$ be a prime minimal over $\operatorname{ker}(\mu)$; then $\mathfrak{q}\in\operatorname{Ass}R$, and localizing at $\mathfrak{q}$ yields $\operatorname{ker}(\mu)_{\mathfrak{q}}\neq 0$, so $\mu_{\mathfrak{q}}=0$, a contradiction. Hence $\operatorname{ker}(\mu)=0$, i.e. $\operatorname{ann}_{R}(x)=0$. Thus $x$ is a torsion-free element of $\omega_{R}$ of degree b=a(R)b=-a(R), as required. ∎
@@
Assume that $R$ is not Gorenstein and $\dim(R)>0$. Consider the following conditions:
@@
    $\omega_{R}]_{-a_{R}}$ contains a torsion-free element of $\omega_{R}$ (e.g., $R$ is a domain), $r_{0}(R)\geq\text{codim}(R)$ and $\operatorname{tr}(\omega_{R})]_{\text{indeg}({\mathfrak{m}})}$ contains a non-zerodivisor;
@@
    $R$ is a level Teter ring with $r(R)=\text{codim}(R)$;
@@
    $R$ is Teter.
Then $(1)\Rightarrow(2)\Rightarrow(3)\Rightarrow(4)$ holds. Moreover, if $R$ is standard graded, all the conditions are equivalent; in particular, if any one of them holds, then $r(R)=r_{0}(R)=\text{codim}(R)$.
@@ 4. Relations between the nearly Gorenstein and Teter properties
In this section, as an application of Section  3, we mainly investigate the relationship between nearly Gorenstein rings and Teter rings. In particular, we show that a certain known class of (non-Gorenstein) nearly Gorenstein standard graded rings is Teter (Section  4), and we clarify the relationship between nearly Gorenstein rings and the Cohen–Macaulay type (Section  4). Throughout this section, we retain Section  1 and use the notation of Section  2. In addition, we assume throughout that $(R,\mathfrak{m})$ is a positively graded Cohen–Macaulay ring.
@@
It is well known that if $R$ is standard graded and has minimal multiplicity, then it is a level ring and $r_{0}(R)=h_{1}={\text{codim}}(R)$ (see [eisenbud1984linear, Section 4]). Hence this follows from Section  3. ∎
@@
As mentioned above, in dimension at least two a ring with minimal multiplicity need not be Teter. To see this, let $n\geq 4$ and let $X=(x_{i,j})_{1\leq i\leq 2,\;1\leq j\leq n}$ be a $2\times n$ generic matrix. Consider the generic determinantal ring $R=\mathbb{k}[X]/I_{2}(X)$ defined by the $2\times 2$ minors of $X$, where $\mathbb{k}$ is a field. Then $R$ is a Cohen–Macaulay standard graded domain; moreover, it is the homogeneous coordinate ring of a rational normal scroll and has minimal multiplicity (see [bruns1998cohen, eisenbud1984linear]). On the other hand, by [ficarra2024canonical!, Theorem 1.1] we have $\operatorname{tr}_{R}(\omega_{R})=I_{1}(X)^{n-2}R={\mathfrak{m}}_{R}^{n-2}$, and in particular $\operatorname{tr}_{R}(\omega_{R})]_{1}=(0)$, so it contains no non-zerodivisor of $R$. Thus, by Section  4 $R$ is not Teter.
We recall that for standard graded Teter rings, the difference between the Cohen–Macaulay type and the number of minimal generators of the canonical module in the smallest degree is always 0. In what follows, we first show that several known classes of nearly Gorenstein rings are Teter. Moreover, we show that there exist semi-standard graded Teter rings for which the above difference can take an arbitrary value.
@@
Then $R$ is Teter. Moreover, for any positive integer $a>0$, there exists a nearly Gorenstein (non-standard graded) semi-standard graded Teter ring $R$ satisfying $(7)$ such that $r(R)-r_{0}(R)=a$.
@@
(4): This follows from Section  4.
(5): By [miyashita2025canonical, Theorem A], $R$ is isomorphic to the Stanley–Reisner ring corresponding to a path of length at least $3$, and in particular has minimal multiplicity. Hence $R$ satisfies (4), and therefore is Teter.
@@
(7): By [miyashita2024comparing, Proposition 5.4 and Theorem 6.3], $R$ is non-level and is a nearly Gorenstein semi-standard graded ring in the sense of [goto2015almost]. Hence, by [miyashita2024comparing, Theorem 6.3], we may write $R=R_{0}[S]$, where
@@
where $\omega_{S}=\big\langle\{\,(1-2i,\,-1+2i):k-1-a\leq i\leq k-1\,\}\cup\{\,(2i,\,2a+2-i):1\leq i\leq a\,\}\big\rangle$. Thus, for the ideal $I=t^{(2k-3,\,2a+5-2k)}\omega_{R}\cong\omega_{R}$, since $I\subseteq R$ and $\operatorname{codim}(R/I)=2$, the ring $R$ is Teter. Moreover, since $r_{0}(R)=a+1$ and $r(R)=2a+1$, we have $r(R)-r_{0}(R)=a$. ∎
Next, we investigate the relationship between the trace ideal of the canonical module and the Cohen–Macaulay type.
@@
    If $r_{0}(R)\geq\text{codim}(R)$, then it is level and $r(R)=\text{codim}(R)$.
@@
(1): This follows from the implication $(1)\Rightarrow(3)$ in Theorem 3.13.
(2): This follows from the implication $(2)\Rightarrow(3)$ in Theorem 3.13. ∎
@@
As a consequence, we obtain the following corollary concerning the Cohen–Macaulay type of nearly Gorenstein rings.
@@
Assume $R$ is non-Gorenstein nearly Gorenstein, and that $R_{\text{indeg}({\mathfrak{m}}_{R})}$ contains a non-zerodivisor of $R$ (e.g., when $R$ is a domain or a semi-standard graded ring with $\dim R>0$). Then the following hold:
@@
    If $r_{0}(R)\geq\text{codim}(R)$, then $R$ is level and $r(R)=\text{codim}(R)$;
@@
    If $R$ is standard graded, $\text{codim}(R)=3$, $\dim(R)\geq 2$ and $r_{0}(R)\neq 2$, then $R$ is level and $r(R)=3$;

### v2

@@ 2.4. Numerical semigroup rings
$\operatorname{PF}(H):=\{h\in H\,:\,\alpha+h\in H\,\;{\it for\;all}\;\,h\in H\},$
$\operatorname{Ap}(H,a_{1})=\{h\in H:h-a_{1}\in H\}.$
@@
and we have $\operatorname{r}(R_{H})=\#(\operatorname{PF}(H))=:\operatorname{type}(H)$.
Several classes of numerical semigroup rings are characterized using $\operatorname{PF}(H)$.
Remark 2.15.
Let $H$ be a numerical semigroup. We define the following classes:
@@
    Suppose that $\operatorname{F}(H)$ is even. We say that $H$ is pseudo-symmetric if $\operatorname{PF}(H)=\left\{\frac{\operatorname{F}(H)}{2},\operatorname{F}(H)\right\}$.
@@
$\Omega_{H}^{+}=\{m\in\mathbb{N}:t^{m}\in\omega_{R_{H}}\},$
@@
$\Omega_{H}^{-}=\{m\in\mathbb{N}:m-\alpha_{i}\in H,for\,all\,i\in[n]\}.$
According to [herzog2019trace, Lemma1.1], the canonical trace ideal $\operatorname{tr}_{R_{H}}(\omega_{R_{H}})$ of $R_{H}$ is obtained by $\operatorname{tr}_{R_{H}}(\omega_{R_{H}})=\omega_{R_{H}}\omega_{R_{H}}^{-1}$. Thus, the set $\operatorname{tr}_{R_{H}}(H)$ corresponding to the generators of $\operatorname{tr}_{R_{H}}(\omega_{R_{H}})$ over $\mathbb{k}$ is given by $\operatorname{tr}(H)=\Omega_{H}^{+}+\Omega_{H}^{-}$.
@@ 3. Graded Teter rings and canonical traces
    Assume that $\dim(R)=0$ and $R$ is not Gorenstein, then the following are equivalent:
@@
$\varphi(\omega_{R})=\mathfrak{m}\quad\text{or}\quad\varphi\text{ is injective and }\operatorname{embdim}\!\bigl(R/\varphi(\omega_{R})\bigr)\leq\dim(R).$
@@
(1) $\Rightarrow$ (4): Choose $\varphi:\omega_{R}\to R$ as the composite of an isomorphism $\omega_{R}\cong I$ with the inclusion $I\hookrightarrow R$. Then $\operatorname{embdim}\!\bigl(R/\varphi(\omega_{R})\bigr)=\operatorname{embdim}(R/I)\leq\dim(R)$, and hence $R$ is Teter.
(4) $\Rightarrow$ (1): By (a) and definition, there exists a graded $R$-homomorphism $\varphi:\omega_{R}\to R$ such that $\varphi$ is injective and $\operatorname{embdim}\bigl(R/\varphi(\omega_{R})\bigr)\leq\dim(R)$. Set $I=\varphi(\omega_{R})$. As $R$ is not Gorenstein, we have $I\neq R$; consequently $\omega_{R}\cong I$ and $\operatorname{embdim}(R/I)\leq\dim(R)$, establishing (1).
@@
When $\dim(R)=0$, it follows from Section  3 (a) that $R$ is Teter in the sense of [Teter1974SocleFactor]. In this case, [elias2017teter, Theorem 3.4] yields $r(R)=\operatorname{embdim}(R)=\text{codim}(R)$. Hence it suffices to treat the case $\dim(R)>0$. By Section  3 (b), the localization $R_{{\mathfrak{m}}}$ is a local Teter ring. Therefore, by [puthenpurakal2025higher, Proposition 5.15] we have $r(R_{{\mathfrak{m}}})=\operatorname{embdim}(R_{{\mathfrak{m}}})-\dim(R_{{\mathfrak{m}}})$. Since $r(R_{{\mathfrak{m}}})=r(R)$, $\operatorname{embdim}(R_{{\mathfrak{m}}})=\operatorname{embdim}(R)$, $\dim(R_{{\mathfrak{m}}})=\dim(R)$ (see [bruns1998cohen, Proposition 1.5.15]), we conclude that $r(R)=\operatorname{embdim}(R)-\dim(R)=\text{codim}(R)$. ∎
Recall that $r_{\operatorname{min}}(R):=\dim_{R_{0}}([\omega_{R}]_{-a_{R}})$.
@@
Assume that $\omega_{R}]_{-a_{R}}$ contains an $R$-torsion free element of $\omega_{R}$. If $\operatorname{tr}(\omega_{R})]_{\text{indeg}_{R}({\mathfrak{m}})}$ contains a non-zero divisor, then there is a graded ideal $I$ of $R$ such that $\omega_{R}\cong I$ as a graded $R$-module and $r_{\operatorname{min}}(R)\leq\operatorname{embdim}(R)-\operatorname{embdim}(R/I).$
@@
By § 2.3, there exists a graded ideal $I\subsetneq R$ such that $\omega_{R}\cong I$. Then, $\operatorname{embdim}(R/I)\leq\operatorname{embdim}(R)-r_{\operatorname{min}}(R)$. This yields the desired inequality. ∎
@@
Assume that $r_{\operatorname{min}}(R)\geq\text{codim}(R)$, $\omega_{R}]_{-a_{R}}$ contains an $R$-torsion free element of $\omega_{R}$, and that $\operatorname{tr}(\omega_{R})]_{\text{indeg}_{R}({\mathfrak{m}})}$ contains a non-zero divisor, then $R$ is Teter.
@@
$\operatorname{embdim}(R)-\dim(R)=\text{codim}(R)\leq r_{\operatorname{min}}(R)\leq\operatorname{embdim}(R)-\operatorname{embdim}(R/I).$
@@
Since $R$ is Teter, by Section  3 there exists a graded ideal $I\subset R$ such that $\omega_{R}\cong I$ and $\dim_{R_{0}}(I_{1})\geq\text{codim}(R)$. In this case $I_{1}\neq 0$, and since $r_{\operatorname{min}}(R)=\dim_{R_{0}}(I_{1})\geq\text{codim}(R)$, we have $r_{\operatorname{min}}(R)\geq\text{codim}(R)$. Therefore, by Section  3 we obtain $\text{codim}(R)\leq r_{\operatorname{min}}(R)\leq r(R)=\text{codim}(R)$. Hence $r(R)=r_{\operatorname{min}}(R)$, so $R$ is level. Because $R$ is Teter with $\dim(R)>0$, it is in particular generically Gorenstein by Section  3 (c). Since $R$ has positive Krull dimension, is level, and is generically Gorenstein, the same argument as in the proof of [bruns1998cohen, Theorem 4.4.9] shows that $I_{1}\ (\cong[\omega_{R}]_{-a_{R}})$ contains a non-zerodivisor of $R$. Consequently $\operatorname{tr}_{R}(\omega_{R})]_{1}=[\operatorname{tr}_{R}(I)]_{1}$ contains $I_{1}$ and, in particular, contains a non-zerodivisor of $R$. ∎
@@
Let $R$ be a Cohen–Macaulay standard graded Teter ring. By Section  3, if $\dim(R)>0$ then $R$ is level, so it suffices to consider the case $\dim(R)=0$. Although the fact that an Artinian graded Teter ring is level is probably well known to experts, we include a proof for completeness. In this case there exists a standard graded Artin Gorenstein ring $G$ such that $R\cong G/\operatorname{Soc}(G)$ and $\operatorname{embdim}(R)=\operatorname{embdim}(G)$, where $\operatorname{Soc}(G):=\{a\in G:a{\mathfrak{m}}_{G}=0\}$. Let $s$ be the socle degree of $G$ and put $n=\operatorname{embdim}(R)$. Then the $h$-vectors of $G$ and $R$ can be written as $h(G)=(h_{0},\dots,h_{s(G)})$ and $h(R)=(h_{0},\dots,h_{s(G)-1})$. Since $G$ is Gorenstein, its $h$-vector is symmetric (see [stanley1978hilbert, Theorem 4.1]), hence $h_{s(G)-1}=h_{1}=\dim_{\mathbb{k}}R_{1}=n$. Therefore $R$ is level by [elias2017teter, Theorem 3.12]. ∎
@@
    Section  3 fails in general in the semi-standard graded setting, and there exist examples in which the difference between $r(R)$ and $r_{\operatorname{min}}(R)$ can take any prescribed value (see Section  4).
@@
    In Section  3 (c), (4) $\Rightarrow$ (3) does not hold in general. Let $A=\mathbb{k}[x,y]/(x^{3},y^{4},xy)$. Note that $B=A_{{\mathfrak{m}}}$ is an Artinian Teter local ring in the sense of [Teter1974SocleFactor] (see [gasanova2022rings, Example 4.3 (2)]). Set $R=A[z]$ and regard $R$ as a standard graded ring by declaring $\operatorname{deg}(x)=\operatorname{deg}(y)=\operatorname{deg}(z)=1$. By [puthenpurakal2025higher, Example 5.15], we have $R_{{\mathfrak{m}}_{R}}\cong B[z]_{{\mathfrak{m}}_{B}B[z]+zB[z]}$ is a Teter local ring in the sense of [puthenpurakal2025higher]. However, since $R$ is not level, $R$ is not Teter by Section  3.
Remark 3.12.
@@
While the statement above differs from [bruns1998cohen, Theorem 4.4.9], the argument in its proof yields the claim needed here. Moreover, the standard graded hypothesis is not used and can be omitted verbatim. ∎
@@
Assume that $R$ is not Gorenstein. Consider the following conditions:
@@
    $\omega_{R}]_{-a_{R}}$ contains a torsion-free element of $\omega_{R}$ (e.g., $R$ is a domain), $r_{\operatorname{min}}(R)\geq\text{codim}(R)$ and $\operatorname{tr}(\omega_{R})]_{\text{indeg}({\mathfrak{m}})}$ contains a non-zerodivisor;
@@
    $R$ is a Teter and level ring with $r(R)=\text{codim}(R)$ and $\dim(R)>0$;
@@
    $R$ is Teter with $\dim(R)>0$.
Then $(1)\Rightarrow(2)\Rightarrow(3)\Rightarrow(4)$ holds. Moreover, if $R$ is standard graded, all the conditions are equivalent; in particular, if any one of them holds, then $r(R)=r_{\operatorname{min}}(R)=\text{codim}(R)$.
@@ 4. Relations between the nearly Gorenstein and Teter properties
In this section, as an application of Section  3, we mainly investigate the relationship between nearly Gorenstein rings and Teter rings. In particular, we show that various known nearly Gorenstein rings are Teter (Section  4), and we clarify the relationship between nearly Gorenstein rings and the their type (Section  4). Throughout this section, we retain Section  1 and use the notation of Section  2. In addition, we assume throughout that $(R,\mathfrak{m})$ is a positively graded Cohen–Macaulay ring.
@@
It is well known that if $R$ is standard graded and has minimal multiplicity, then it is a level ring and $r_{\operatorname{min}}(R)=h_{1}={\text{codim}}(R)$ (see [eisenbud1984linear, Section 4]). Hence this follows from Section  3. ∎
@@
As mentioned above, in dimension at least two a ring with minimal multiplicity need not be Teter. To see this, let $n\geq 4$ and let $X=(x_{i,j})_{1\leq i\leq 2,\;1\leq j\leq n}$ be a $2\times n$ generic matrix. Consider the generic determinantal ring $R=\mathbb{k}[X]/I_{2}(X)$ defined by the $2\times 2$ minors of $X$. Then $R$ is a Cohen–Macaulay standard graded domain; moreover, it is the homogeneous coordinate ring of a rational normal scroll and has minimal multiplicity (see [bruns1998cohen, eisenbud1984linear]). On the other hand, by [ficarra2024canonical!, Theorem 1.1] we have $\operatorname{tr}_{R}(\omega_{R})=I_{1}(X)^{n-2}R={\mathfrak{m}}_{R}^{n-2}$, and in particular $\operatorname{tr}_{R}(\omega_{R})]_{1}=(0)$, so it contains no non-zerodivisor of $R$. Thus, by Section  4 $R$ is not Teter.
We recall that for standard graded Teter rings, the difference between the type and the number of minimal generators of the canonical module in the smallest degree is always 0. In what follows, we first show that several known classes of nearly Gorenstein rings are Teter. Moreover, we show that there exist semi-standard graded Teter rings for which the above difference can take an arbitrary value.
@@
Then $R$ is Teter. Moreover, for any positive integer $a>0$, there exists a nearly Gorenstein (non-standard graded) semi-standard graded Teter ring $R$ satisfying $(7)$ such that $r(R)-r_{\operatorname{min}}(R)=a$.
@@
(4) and (5): (4) follows from Section  4. By [miyashita2025canonical, Theorem A], $R$ is isomorphic to the Stanley–Reisner ring corresponding to a path of length at least $3$, and in particular has minimal multiplicity. Hence $R$ satisfies (4), and therefore is Teter.
@@
(7): By [miyashita2024comparing, Proposition 5.4 and Theorem 6.3], $R$ is non-level and is a nearly Gorenstein semi-standard graded ring in the sense of [goto2015almost]. Hence, by [miyashita2024comparing, Theorem 6.3], we may write $R=\mathbb{k}[S]$, where
@@
where $\omega_{S}=\big\langle\{\,(1-2i,\,-1+2i):k-1-a\leq i\leq k-1\,\}\cup\{\,(2i,\,2a+2-i):1\leq i\leq a\,\}\big\rangle$. Thus, for the ideal $I=t^{(2k-3,\,2a+5-2k)}\omega_{R}\cong\omega_{R}$, since $I\subset R$ and $\operatorname{codim}(R/I)=2$, the ring $R$ is Teter. Moreover, since $r_{\operatorname{min}}(R)=a+1$ and $r(R)=2a+1$, we have $r(R)-r_{\operatorname{min}}(R)=a$. ∎
Next, we investigate the relationship between the trace ideal of the canonical module and the type.
@@
    If $r_{\operatorname{min}}(R)\geq\text{codim}(R)$, then it is level and $r(R)=\text{codim}(R)$.
@@
(1) follows from the implication $(1)\Rightarrow(3)$ in Theorem 3.13. (2) follows from the implication $(2)\Rightarrow(3)$ in Theorem 3.13. ∎
@@
As a consequence, we obtain the following corollary concerning the type of nearly Gorenstein rings.
@@
Assume $R$ is non-Gorenstein nearly Gorenstein, and that $R_{\text{indeg}({\mathfrak{m}}_{R})}$ contains a non-zero divisor of $R$ (e.g., when $R$ is a domain or a semi-standard graded ring with $\dim(R)>0$). Then the following hold:
@@
    If $r_{\operatorname{min}}(R)\geq\text{codim}(R)$, then $R$ is level and $r(R)=\text{codim}(R)$;
@@
    If $R$ is standard graded, $\text{codim}(R)=3$, $\dim(R)\geq 2$ and $r_{\operatorname{min}}(R)\neq 2$, then $R$ is level and $r(R)=3$;